# Optimization-based control by interconnection of nonlinear port-Hamiltonian systems

### v1

@@ 1. Introduction
However, for small sampling times and in view of increasing complexity of real-world applications such as problems governed by partial differential equations, it can be challenging to compute the optimal solution in a short time. As a consequence, there is need for MPC approaches that utilize suboptimal solutions as feedback, obtained e.g. by an iterative optimization algorithm. In these methods, the solution of the optimal control problem on the bottom right of Figure 1 is replaced by an inexact solution given, e.g., by the truncation of an iterative optimization method. In recent decades, various suboptimal MPC schemes for finite-dimensional nonlinear control systems have been proposed and analyzed [8, 29, 11, 20].
@@
To address this issue, we propose a control by interconnection approach in the spirit of suboptimal MPC by formulating a continuous-time primal-dual gradient method for nonlinear control-constrained optimal control problems as a port-Hamiltonian system. To this end, we interconnect the dynamics of gradient-type optimization algorithm via an MPC-type feedback with a port-Hamiltonian plant in a structure-preserving-way, leading to a dissipative even asymptotically stable closed-loop system depicted in Figure 2. We provide stability results for primal-dual gradient dynamics and prove convergence of the interconnected system in function space. This is particularly central as the convergence does not rely on any particular sampling or discretization scheme, as long as it preserves the port-Hamiltonian structure. Last, the proposed analysis in function space paves the way towards suboptimal MPC for partial differential equations, a topic that has been largely unexplored so far. In this context we mention related works. Finite-dimensional saddle-point dynamics occurring in primal-dual gradient methods have been analyzed in [7, 6] and the port-Hamiltonian structure has been observed in [22] for stationary finite-dimensional problems. In [18, 26], the authors suggest an MPC-type control-by interconnection of port-Hamiltonian systems for discrete-time finite-dimensional optimal control problems.
@@
Our main contribution, situated at the interface of control-by-interconnection and structure-exploiting optimal control for pHs, and the structure of this work are as follows. First, in Section 2, we introduce a class of infinite-dimensional port-Hamiltonian systems subject to set-valued nonlinear operators that is suitable to model primal-dual gradient dynamics for continuous-time control-constrained optimal control problems. The second contribution is centered around formulation (Section 3) and analysis (Section 4) of the primal-dual gradient dynamics for continuous-time control-constrained problems. In Theorem 4.7, we establish weak convergence of the primal–dual gradient dynamics to the unique optimal point of the underlying optimization problem. The proof is based on an infinite-dimensional LaSalle invariance principle leveraging the convexity of the problem and a detectability-like assumption of the cost function. The corresponding result for control-constrained problems is given in Theorem 4.8. Third, we prove stability of a suboptimal MPC-type control-by-interconnection method for unconstrained stabilization of pHs in Section 5, namely in Theorems 5.2 and 5.3 based on either coervicity or system-theoretic properties such as detectability and controllability using again a LaSalle-type argument. An extension to control-constrained stabilization is provided in Theorem 5.6 under an observability condition. We conclude our work by means of a numerical experiment in Section 6. The proofs of two theorems related to our LaSalle-type result are moved to Appendix 8.1. Some important notions and results from convex analysis are recalled in Appendix 8.2.
@@ 2.2. Structure-preserving interconnection
for all $x,y\in D$, see [2, Prop. 4.4].
@@
Let $(M_{1},B_{1})$ and $(M_{2},B_{2})$ be two (maximally) monotone pH systems in the Hilbert spaces $X_{1}$ and $X_{2}$, with inputs and outputs $u_{i},y_{i}\in U_{i}$, $i=1,2$. Let $H$ be another Hilbert space, $\mathcal{E}:H\to H$ be firmly nonexpansive and $\mathcal{F}_{i}\in L(U_{i},H)$, $i=1,2$. Define the coupling
@@
defines a (maximally) monotone pH system, where
@@
We decompose $M=N+K$ where $N=\left[\begin{smallmatrix}M_{1}\\ M_{2}\end{smallmatrix}\right]$. Since $N$ is maximally monotone it is, again due to the sum rule for maximally monotone operators, enough to prove that $K$ is monotone and $D(K)=X_{1}\times X_{2}$. To this end, consider for $x_{i}\in D(M_{i})$, $i=1,2$:
@@
where the first inequality follows by firm nonexpansiveness of $\mathcal{E}$, [2, Prop. 4.2]. Hence, $K$ is monotone, and its maximality is a consequence of $\mathcal{E}$ and $\mathcal{F}_{i}$ being everywhere defined. ∎
@@ 3. Monotonicity in optimal control
where the last inequality follows as $F$ is bounded. Since $F$ is a bounded subset of a the Hilbert space $U$, $F$ is relatively weakly sequentially compact, [23, Thm. 2.10]. Therefore, $(f_{n})_{n\in\mathbb{N}}$ admits a weakly convergent subsequence $(f_{n_{k}})_{k\in\mathbb{N}}\subset F$ with
@@
as $k\to\infty$. But $F$ is closed and convex and with that, $F$ is weakly sequentially closed, [23, Thm. 2.11], which proves that $f^{\star}\in F$. Hence,
@@ 4.1. Stability of unconstrained optimizer dynamics
The proof of the following theorem can be found in [21, Thm. 2.3].
@@
by [25, Lem. A.9]. Since $(T_{\widetilde{M_{\mathrm{opt}}^{0}}}(t))_{t\geq 0}$ is a semigroup of nonexpansive mappings, we obtain
@@ 4.2. Stability of constrained optimizer dynamics
which holds for all $t\geq 0$. Applying the Gronwall Lemma from [12, Lemma 4.1] leads to
@@ 5.2. Inequality constrained problems
By Proposition 2.4, the interconnection (5.11) leads to an autonomous maximally monotone pH system (5.12) on $\mathcal{V}_{\mathrm{r}}$.
@@ 6. Numerical examples
for some choice $\overline{u}=-\underline{u}\in\mathbb{R}_{+}$. We approach the two evolution equations (5.6) (unconstrained case) and (5.12) (constrained case) numerically. To this end, the differential operators $\mathcal{C}$ and $\mathcal{C}^{*}$ are approximated by the implicit Euler rule and the resulting finite-dimensional system is solved using an explicit Runge-Kutta method with a very fine resolution. The norm of the plant solution over time is shown in Figure 3. The feedback controls that are applied to the plant can be seen in Figure 4. It can be observed that the control in the constrained case is pointwisely contained in the feasible set $F$. The norm of the solution of the full coupled system over time is illustrated in 5. Observe that, due to monotonicity of the governing operators, the map $t\mapsto\|w(t)\|^{2}$ is monotonically nonincreasing. Indeed, the numerical approximation of the fully coupled system is observed to be asymptotically stable, indicating convergence of the Lagrange multipliers $(\lambda,\lambda_{0},\mu)(t)$ to the unique optimal point of the underlying optimality system (3.8).

### v2

@@ 1. Introduction
However, for small sampling times and in view of increasing complexity of real-world applications such as problems governed by partial differential equations, it can be challenging to compute the optimal solution in a short time. As a consequence, there is need for MPC approaches that utilize suboptimal solutions as feedback, obtained e.g. by an iterative optimization algorithm. In these methods, the solution of the optimal control problem on the bottom right of Figure 1 is replaced by an inexact solution given, e.g., by the truncation of an iterative optimization method. In recent decades, various suboptimal MPC schemes for finite-dimensional nonlinear control systems have been proposed and analyzed, see [18, Section 8.9.2] and [12, Section 10.6]. In the seminal paper [19], the authors leverage non-optimal solutions for MPC with terminal ingredients, as long as a suitable Lyapunov decrease condition is satisfied. Another line of research is embedded in the context of continuation methods [18, Section 8.9.2] or initial-value embeddings [12, p. 325ff] for optimization problems. One of the most prominent representative is the real-time iteration [8]. There, one pursues a Newton step for the optimal control problem and it may be shown that, under suitable assumptions on the sampling time, the method yields asymptotic stability [8], see also the recent work [29] for a Lyapunov-based analysis of the coupled optimizer-plant dynamics. Third, and related to our suggested approach, we mention instant MPC (see [28] for the case of linear dynamics), in which only one primal-dual gradient step is performed in the feedback computation. Despite the large body of literature regarding suboptimal MPC, the majority of the works provide feedback schemes in the absence of inequality constraints. Further, none of the works consider a function-space setting, which already is necessary for time-continuous optimal control problems as shown in this work and in particular is crucial to analyze suboptimal MPC for PDEs in future work, a topic which is widely unexplored in the literature.
@@
To address this gap, we propose a control by interconnection approach in the spirit of suboptimal MPC by formulating a continuous-time primal-dual gradient method for nonlinear control-constrained optimal control problems as a port-Hamiltonian system. To this end, we interconnect the dynamics of gradient-type optimization algorithm via an MPC-type feedback with a port-Hamiltonian plant in a structure-preserving way, leading to a dissipative even asymptotically stable closed-loop system depicted in Figure 2. We provide stability results for primal-dual gradient dynamics and prove convergence of the interconnected system in function space. This is particularly crucial, as the convergence does not rely on any particular sampling or discretization scheme, as long as it preserves the port-Hamiltonian structure. Finally, the proposed analysis in function space paves the way toward suboptimal MPC for partial differential equations, a topic that has remained largely unexplored. In this context, we mention related works. Finite-dimensional saddle-point dynamics occurring in primal-dual gradient methods have been analyzed in [7, 6] and the port-Hamiltonian structure has been observed in [21] for stationary finite-dimensional problems. In [17, 25], the authors suggest an MPC-type control-by interconnection of port-Hamiltonian systems for discrete-time finite-dimensional optimal control problems.
@@
Contribution and outline. Our main contribution, situated at the interface of control-by-interconnection and structure-exploiting optimal control for pHs, and the structure of this work are as follows. First, in Section 2, we introduce a class of infinite-dimensional port-Hamiltonian systems subject to set-valued nonlinear operators that is suitable to model primal-dual gradient dynamics for continuous-time control-constrained optimal control problems. The second contribution is centered around formulation (Section 3) and analysis (Section 4) of the primal-dual gradient dynamics for continuous-time control-constrained problems. In Theorem 4.7, we establish weak convergence of the primal–dual gradient dynamics to the unique optimal point of the underlying optimization problem. The proof is based on an infinite-dimensional LaSalle invariance principle leveraging the convexity of the problem and a detectability-like assumption of the cost function. The corresponding result for control-constrained problems is given in Theorem 4.8. Third, we prove stability of a suboptimal MPC-type control-by-interconnection method for unconstrained stabilization of pHs in Section 5, namely in Theorems 5.2 and 5.3 based on either coercivity or system-theoretic properties such as detectability and controllability using again a LaSalle-type argument. An extension to control-constrained stabilization is provided in Theorem 5.6 under an observability condition. We conclude our work by means of a numerical experiment in Section 6.
Code availability. The code for all numerical experiments is provided in the repository
https://github.com/preustertill/suboptimal_MPC.git.
@@ 2.2. Structure-preserving interconnection
for all $x,y\in D$, see [2, Prop. 4.4]. We call a set-valued operator $M:X\rightrightarrows X$ relative monotone with respect to 0 if (2.1) holds with $x_{2}=y_{2}=0$ and the corresponding system (mpHS) is called relative monotone pHs.
@@
Let $(M_{1},B_{1})$ and $(M_{2},B_{2})$ be two relative monotone pH systems in the Hilbert spaces $X_{1}$ and $X_{2}$, with inputs and outputs $u_{i},y_{i}\in U_{i}$, $i=1,2$. Let $H$ be another Hilbert space, $\mathcal{E}:H\to H$ be firmly nonexpansive and $\mathcal{F}_{i}\in L(U_{i},H)$, $i=1,2$. Define the coupling
@@
defines a relative monotone pH system, where
@@
We decompose $M=N+K$ where $N=\left[\begin{smallmatrix}M_{1}\\ M_{2}\end{smallmatrix}\right]$ is clearly relative monotone. Hence, it suffices to prove that $K$ is relative monotone. Thus for $x_{i}\in D(M_{i})$, $i=1,2$,
@@
where the first inequality follows by firm nonexpansiveness of $\mathcal{E}$, [2, Prop. 4.2]. ∎
@@ 3. Monotonicity in optimal control
where the last inequality follows as $F$ is bounded. Since $F$ is a bounded subset of a the Hilbert space $U$, $F$ is relatively weakly sequentially compact, [22, Thm. 2.10]. Therefore, $(f_{n})_{n\in\mathbb{N}}$ admits a weakly convergent subsequence $(f_{n_{k}})_{k\in\mathbb{N}}\subset F$ with
@@
as $k\to\infty$. But $F$ is closed and convex and with that, $F$ is weakly sequentially closed, [22, Thm. 2.11], which proves that $f^{\star}\in F$. Hence,
@@ 4.1. Stability of unconstrained optimizer dynamics
The proof of the following theorem can be found in [20, Thm. 2.3].
@@
by [24, Lem. A.9]. Since $(T_{\widetilde{M_{\mathrm{opt}}^{0}}}(t))_{t\geq 0}$ is a semigroup of nonexpansive mappings, we obtain
@@ 4.2. Stability of constrained optimizer dynamics
which holds for all $t\geq 0$. Applying the Gronwall Lemma from [11, Lemma 4.1] leads to
@@ 5.2. Inequality constrained problems
By Proposition 2.4, the interconnection (5.11) leads to an autonomous relative monotone pH system (5.12) on $\mathcal{V}_{\mathrm{r}}$.
@@ 6. Numerical examples
for some choice $\overline{u}=-\underline{u}\in\mathbb{R}_{+}$. We approach the two evolution equations (5.6) (unconstrained case) and (5.12) (constrained case) numerically. To this end, the differential operators $\mathcal{C}$ and $\mathcal{C}^{*}$ are approximated by the implicit Euler rule and the resulting finite-dimensional system is solved using an explicit Runge-Kutta method with a very fine resolution. The norm of the plant solution over time is shown in Figure 3. The feedback controls that are applied to the plant can be seen in Figure 4. It can be observed that the control in the constrained case is pointwisely contained in the feasible set $F$. The norm of the solution of the full coupled system over time is illustrated in 5. Observe that, due to monotonicity of the governing operators, the map $t\mapsto\|w(t)\|^{2}$ is monotonically nonincreasing. Indeed, the numerical approximation of the fully coupled system is observed to be asymptotically stable, indicating convergence of the Lagrange multipliers $(\lambda,\lambda_{0})(t)$ to the unique optimal point of the underlying optimality system (3.8).